You just paid $329,000 for an annuity that will pay you and your heirs $11,900 a year forever. What rate of return are you earning on this policy?

a. 3.77 percent

b. 3.96 percent

c. 3.62 percent

d. 3.37 percent

e. 3.68 percent
Answer #1

Answer is (c) 3.62%

This is a question of Present value of perpetuity

PV of perpetuity=d/r

where d = periodic return

r = rate of return

rate of return(r) shall be = d/pv of perpetuity

=11900/329000

= .03617 or 3.62%(approx).

I hope you understood the concept.

Please do not forget to upvote my solution if you like the same.

Feel free to ask any query via comments.

Good Luck!